voidedits1:

What is the force of gravity between Earth (6.0 × 1024 kilograms) and Venus (4.88 × 1024 kilograms)? The distance between the two planets is about 3.8 × 1010 meters. (The value of G is 6.673 × 10-11 newton meter2/kilogram2. The mass of Earth is 5.98 × 1024 kilograms.)

IamAku:

The formula is F=Gm1m2/d^2. m1=mass of earth m2=mass of venus d=distance between earth nd venus I hope this doesn't count as a direct answer :')

voidedits1:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @IamAku The formula is F=Gm1m2/d^2. m1=mass of earth m2=mass of venus d=distance between earth nd venus I hope this doesn't count as a direct answer :') \(\color{#0cbb34}{\text{End of Quote}}\) :) thank you, I appreciate your help

IamAku:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @voidedits1 \(\color{#0cbb34}{\text{Originally Posted by}}\) @IamAku The formula is F=Gm1m2/d^2. m1=mass of earth m2=mass of venus d=distance between earth nd venus I hope this doesn't count as a direct answer :') \(\color{#0cbb34}{\text{End of Quote}}\) :) thank you, I appreciate your help \(\color{#0cbb34}{\text{End of Quote}}\) Ofc